# An Adaptive Algorithm for Bilevel Optimization on Riemannian Manifolds

### v2

@@ 3 Adaptive Riemannian hypergradient descent algorithms
Standard Riemannian bilevel optimization (RBO) approaches [19, 29, 48] determine step sizes for updating the variables using problem-specific parameters such as strong convexity, Lipschitzness, and curvature constants. However, these parameters are frequently impractical to estimate or compute, posing challenges for step size determination. This limitation underscores the need for adaptive RBO algorithms that operate without prior parameter knowledge.
@@ 3.1 Approximate Riemannian hypergradient
[29, 48] The Riemannian hypergradient of $F(x)$ are given by
@@
The aforementioned proposition fundamentally hinges on the implicit function theorem for Riemannian manifolds [27], which necessitates the invertibility of the Hessian of the lower-level objective function $g$ w.r.t. $y$. In this paper, this requirement is satisfied as $g$ is geodesically strongly convex w.r.t. $y$. Consequently, $y^{*}(x)$ is unique and differentiable.
However, the Riemannian hypergradient defined in (3) is computationally challenging to evaluate. First, the exact solution $y^{*}(x)$ of the lower-level objective is not explicitly available mostly, necessitating the use of an approximate solution $\hat{y}$, we employ the adaptive Riemannian gradient descent method to compute this approximation. Second, calculating the Hessian-inverse ${\mathcal{H}}_{y}^{-1}g(x,y^{*}(x))[{\mathcal{G}}_{y}f(x,y^{*}(x))]$ in (3) incurs prohibitive computational costs. To address this, given the approximation $\hat{y}$ of $y^{*}(x)$, since $g$ is geodetic strongly convex, we can approximate the Hessian-inverse-vector product ${\mathcal{H}}_{y}^{-1}g(x,\hat{y})[{\mathcal{G}}_{y}f(x,\hat{y})]$ by solving the linear system ${\mathcal{H}}_{y}g[v](x,\hat{y})={\mathcal{G}}_{y}f(x,\hat{y})$, which is originated from the following quadratic problem:
@@
Denote $\hat{v}$ as the approximate solution of Problem (4). In this paper, we use adaptive gradient descent method (cf. Steps 11-15 of Algorithm 1) or tangent space conjugate gradient method [69, 7] (cf. Step 16) to obtain such a $\hat{v}$. Specifically, given an initialization $v^{0}\in T_{\hat{y}}{\mathcal{M}}_{y}$ and error tolerance $\epsilon_{v}$, the tangent space conjugate gradient algorithm for solving the linear system ${\mathcal{H}}_{y}g[v](x,\hat{y})={\mathcal{G}}_{y}f(x,\hat{y})$ is presented in Algorithm 2 of Appendix C.
Therefore, given the estimations $\hat{y}$ and $\hat{v}$, the approximate Riemannian hypergradient [19, 29, 48] is defined as follows,
@@ 3.2 Adaptive Riemannian hypergradient descent algorithm: AdaRHD
Motivated by the recent adaptive methods [81] for solving general bilevel optimization, we employ the “inverse of cumulative gradient norm” strategy to design the adaptive step sizes [80, 78, 81], i.e., adapting the step sizes based on accumulated Riemannian (hyper)gradient norms.
@@
19:     $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$,   $x_{t+1}={\rm Exp}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}))$.
20:end for
@@ 3.3.1 Definitions and assumptions
Similar to the definition of an $\epsilon$-stationary point in general bilevel optimization [24, 38, 39], the definition of an $\epsilon$-stationary point [19, 29, 48] in Riemannian bilevel optimization is given as follows.
@@
The manifolds ${\mathcal{M}}_{x}$ and ${\mathcal{M}}_{y}$ are complete Riemannian manifolds. Moreover, ${\mathcal{M}}_{y}$ is a Hadamard manifold whose sectional curvature is lower bounded by $\tau<0$. Moreover, for all iterates $y_{t}^{k}$ in the lower-level problem, the curvature constant $\zeta(\tau,d(y_{t}^{k},y^{*}(x_{t})))$ defined in Lemma B.1 is always upper bounded by a constant $\zeta$ for all $t\geq 0$ and $p\geq 0$.
Assumption 3.1 is commonly used in the convex Riemannain optimization, which ensures that the lower-level objective $g$ can be geodesically strongly convex [86, 48] and is essential for the convergence of the lower-level problem [29, 48].
@@
Function $f(x,y)$ is continuously differentiable and $g(x,y)$ is twice continuously differentiable for all $(x,y)\in{\mathcal{U}}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$, and $g(x,y)$ is $\mu$-geodesic strongly convex in $y\in{\mathcal{U}}_{y}$ for any $x\in{\mathcal{U}}_{x}$.
This assumption is prevalent in (Riemannian) bilevel optimization [24, 38, 39, 51, 32, 46, 19, 29, 48]. Under this assumption, the optimal solution $y^{*}(x)$ of the lower-level problem is unique for all $x\in{\mathcal{M}}_{x}$ and differentiable [29, 48], ensuring the existence of the Riemannian hypergradient (3.1).
@@
Function $f(x,y)$ is $l_{f}$-Lipschitz continuous in ${\mathcal{U}}\in{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$. The Riemannian gradients ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$ are $L_{f}$-Lipschitz continuous in ${\mathcal{U}}$. The Riemannian gradients ${\mathcal{G}}_{x}g(x,y)$ and ${\mathcal{G}}_{y}g(x,y)$ are $L_{g}$-Lipschitz continuous in ${\mathcal{U}}$. Further, the Riemannian Hessian ${\mathcal{H}}_{y}g(x,y)$, cross derivatives ${\mathcal{G}}^{2}_{xy}g(x,y)$, ${\mathcal{G}}^{2}_{yx}g(x,y)$ are $\rho$-Lipschitz in ${\mathcal{U}}$.
Assumption 3.3 is a standard condition in the existing literature on (Riemannian) bilevel optimization [24, 38, 39, 29, 48]. These conditions ensure the smoothness of the function $F$ and the Riemannian hypergradient ${\mathcal{G}}F$.
@@
The minimum of $F$ over ${\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$, denote as $F^{*}$, is lower-bounded.
@@ 3.3.2 Convergence results
Lemma 3.1 establishes that the error between the approximate and exact Riemannian hypergradients is bounded by the errors arising from the resolutions of the lower-level problem in (1) and the linear system (4).
Similar to Proposition 2 in [81], we derive upper bounds on the total iteration count required to solve both the lower-level problem in Problem (1) and the linear system (4). However, owing to the geometric structures inherent in Riemannian manifolds, these bounds cannot be directly inferred from [81, Proposition 2] and require a meticulous analysis.
@@
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. Then, for any $0\leq t\leq T$, the iterations $K_{t}$ and $N_{t}$ required in Algorithm 1 satisfy:
@@
Proposition 3.2 provides upper bounds for the total iterations of the resolutions of the lower-level problem in Problem (1) and the linear system 4. If $\epsilon_{y}>0$ is sufficiently small, it holds that $K_{t}=\mathcal{O}(1/\epsilon_{y})$, which matches the standard AdaGrad-Norm for solving strongly convex problems [80]. Additionally, if $\epsilon_{v}>0$ is small enough, we also have $N_{t}=\mathcal{O}(1/\epsilon_{v})$ and $N_{t}=\mathcal{O}(\log(1/\epsilon_{v}))$ when employing gradient descent and conjugate gradient methods, respectively.
@@
Theorem 3.1 establishes that the proposed adaptive algorithm achieves a convergence rate matching standard non-adaptive (Riemannian) bilevel algorithms (e.g., [24, 38, 39, 29, 48]), demonstrating its computational efficiency. A complexity bound for Algorithm 1 is subsequently derived.
@@
Corollary 3.1 demonstrates that the gradient complexity ($G_{g}$) and Hessian-vector product complexity ($HV_{g}$) of AdaRHD-GD surpass those of well-tuned Riemannian bilevel optimization methods [29, 48] by a factor of $1/\epsilon$. This gap originates from the additional iterations needed to guarantee the solution of the lower-level problem in Problem (1) and linear system 4, which is necessitated by the lack of prior knowledge regarding the strong convexity, Lipschitzness, and curvature constants.
@@ 3.4 Extend to retraction mapping
All the iterates of the lower-level problem generated by Algorithm 3 lie in a bounded set belonging to ${\mathcal{M}}_{y}$ that contains the optimal solution, i.e., for all $t\geq 0$ there exists a constant $\bar{D}>0$, such that $d(y_{t}^{k},y^{*}(x_{t}))\leq\bar{D}$ for all $k\geq 0$.
@@
Given $z_{1}\in{\mathcal{U}}\in{\mathcal{M}}$ (here ${\mathcal{M}}$ can be ${\mathcal{M}}_{x}$ or ${\mathcal{M}}_{y}$) and $u\in T_{x}{\mathcal{M}}$, let $z_{2}={\rm Retr}_{z_{1}}(u)$. There exist constants $c_{u}\geq 1,c_{R}\geq 0$ such that $d^{2}(z_{1},z_{2})\leq c_{u}\|u\|_{z_{1}}^{2}$ and $\|{\rm Exp}^{-1}_{z_{1}}(z_{2})-u\|_{z_{1}}\leq c_{R}\|u\|_{z_{1}}^{2}$.
Assumption 3.6 is standard (e.g., [29, Assumption 2] and [41, Assumption 1]) in bounding the error between the exponential mapping and retraction mapping, given that retraction mapping is a first-order approximation to the exponential mapping [29].
@@
Suppose that Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, and 3.6 hold. The sequence $\{x_{t}\}_{t=0}^{T}$ generated by Algorithm 3 staisifes,
@@
Furthermore, the complexity analysis of Algorithm 3 closely aligns with that of Algorithm 1 and is, therefore, omitted.
@@ 4 Experimental results
In this section, adhering to the experimental framework established by [29], we conduct comprehensive experiments to evaluate our algorithm against the RHGD method. Since the RieBO algorithm proposed by [48] shares the same algorithmic framework as RHGD, we categorize them as a single method for comparison. We designate RHGD-50 and RHGD-20 as configurations with maximum iteration limits for solving the lower-level problem in RHGD set to 50 and 20, respectively. For consistency, we exclusively employ Algorithm 3 in this section due to its computational efficiency in practice. Detailed experimental settings and additional experiments are provided in Appendix I.

### v3

@@ 3 Adaptive Riemannian hypergradient descent algorithms
Standard Riemannian bilevel optimization (RBO) approaches (dutta2024riemannian, ; han2024framework, ; li2025riemannian, ) determine step sizes for updating the variables using problem-specific parameters such as strong convexity, Lipschitzness, and curvature constants. However, these parameters are frequently impractical to estimate or compute, posing challenges for step size determination. This limitation underscores the need for adaptive RBO algorithms that operate without prior parameter knowledge.
@@ 3.1 Approximate Riemannian hypergradient
(han2024framework, ; li2025riemannian, ) The Riemannian hypergradient of $F(x)$ are given by
@@
Proposition 3.1 fundamentally hinges on the implicit function theorem for Riemannian manifolds han2023nonconvex , which necessitates the invertibility of the Hessian of the lower-level objective function $g$ w.r.t. $y$. In this paper, this requirement is satisfied as $g$ is geodesically strongly convex w.r.t. $y$. Consequently, $y^{*}(x)$ is unique and differentiable.
[…]
@@
Denote $\hat{v}$ as the approximate solution of Problem (4). In this paper, we use the adaptive gradient descent method (cf. Steps 11-15 of Algorithm 1) or the tangent space conjugate gradient method (trefethen2022numerical, ; boumal2023introduction, ) (cf. Step 16) to obtain such a $\hat{v}$. Specifically, the tangent space conjugate gradient algorithm is presented in Algorithm 2 of Appendix C.
Then, given the estimations $\hat{y}$ and $\hat{v}$, the approximate Riemannian hypergradient (dutta2024riemannian, ; han2024framework, ; li2025riemannian, ) is defined as follows,
@@ 3.2 Adaptive Riemannian hypergradient descent algorithm: AdaRHD
Motivated by the recent adaptive methods (yang2024tuning, ) for solving general bilevel optimization, we employ the “inverse of cumulative (Riemannian) gradient norm” strategy to design the adaptive step sizes (xie2020linear, ; ward2020adagrad, ; yang2024tuning, ), i.e., adapting the step sizes based on accumulated Riemannian (hyper)gradient norms.
@@
19:  $a_{t+1}^{2}=a_{t}^{2}+\|\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}$,
20:  $x_{t+1}={\rm Exp}_{x_{t}}(-\frac{1}{a_{t+1}}\widehat{{\mathcal{G}}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}}))$.
21:end for
@@ 3.3.1 Definitions and assumptions
Similar to the definition of an $\epsilon$-stationary point in general bilevel optimization (ghadimi2018approximation, ; ji2021bilevel, ; ji2022will, ), the definition of an $\epsilon$-stationary point in Riemannian bilevel optimization (dutta2024riemannian, ; han2024framework, ; li2025riemannian, ) is given as follows.
@@
The manifolds ${\mathcal{M}}_{x}$ and ${\mathcal{M}}_{y}$ are complete Riemannian manifolds. Moreover, ${\mathcal{M}}_{y}$ is a Hadamard manifold whose sectional curvature is lower bounded by $\tau<0$. Moreover, for all iterates $y_{t}^{k}$ in the lower-level problem, the curvature constant $\zeta(\tau,d(y_{t}^{k},y^{*}(x_{t})))$ defined in Lemma B.1 is always upper bounded by a constant $\zeta$ for all $t\geq 0$ and $k\geq 0$.
Assumption 3.1 is commonly used in the Riemannian optimization, which ensures that the lower-level objective $g$ can be geodesically strongly convex (zhang2016first, ; li2025riemannian, ) and is essential for the convergence of the lower-level problem (han2024framework, ; li2025riemannian, ).
@@
Function $f(x,y)$ is continuously differentiable and $g(x,y)$ is twice continuously differentiable for all $(x,y)\in{\mathcal{U}}={\mathcal{U}}_{x}\times{\mathcal{U}}_{y}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$, and $g(x,y)$ is $\mu$-geodesic strongly convex in $y\in{\mathcal{U}}_{y}$ for any $x\in{\mathcal{U}}_{x}$.
This assumption is prevalent in (Riemannian) bilevel optimization (ghadimi2018approximation, ; ji2021bilevel, ; ji2022will, ; liu2022bome, ; hong2023two, ; kwon2023fully, ; dutta2024riemannian, ; han2024framework, ; li2025riemannian, ). Under this assumption, the optimal solution $y^{*}(x)$ of the lower-level problem is unique for all $x\in{\mathcal{M}}_{x}$ and differentiable (han2024framework, ; li2025riemannian, ), ensuring the existence of the Riemannian hypergradient (3.1).
@@
Function $f(x,y)$ is $l_{f}$-Lipschitz continuous in ${\mathcal{U}}\subseteq{\mathcal{M}}_{x}\times{\mathcal{M}}_{y}$. The Riemannian gradients ${\mathcal{G}}_{x}f(x,y)$ and ${\mathcal{G}}_{y}f(x,y)$ are $L_{f}$-Lipschitz continuous in $\mathcal{U}$}. The Riemannian gradients ${\mathcal{G}}_{x}g(x,y)$ and ${\mathcal{G}}_{y}g(x,y)$ are $L_{g}$-Lipschitz continuous in $\mathcal{U}$}. Furthermore, the Riemannian Hessian ${\mathcal{H}}_{y}g(x,y)$, cross derivatives ${\mathcal{G}}^{2}_{xy}g(x,y)$, ${\mathcal{G}}^{2}_{yx}g(x,y)$ are $\rho$-Lipschitz in $\mathcal{U}$}.
Assumption 3.3 is a standard condition in the existing literature on (Riemannian) bilevel optimization (ghadimi2018approximation, ; ji2021bilevel, ; ji2022will, ; han2024framework, ; li2025riemannian, ). These conditions ensure the smoothness of the objective function $F$ in Problem (1) and the Riemannian hypergradient ${\mathcal{G}}F$ defined in (3).
@@
The minimum of $F$ over ${\mathcal{M}}_{x}$, denote as $F^{*}$, is lower-bounded.
@@ 3.3.2 Convergence results
Lemma 3.1 establishes that the error between the approximate and exact Riemannian hypergradients is bounded by the errors arising from the resolutions of the lower-level problem in Problem (1) and the linear system (4). Consequently, we can actually integrate the stopping criterion $\|\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})\|_{x_{t}}^{2}\leq 1/T$ into Algorithm 1 to enable early termination. By Lemma 3.1, this criterion ensures that $\|{\mathcal{G}}F(x_{t})\|_{x_{t}}^{2}\leq\mathcal{O}({1}/{T})$, which implies that $x_{t}$ is an $\mathcal{O}(1/T)$-stationary point of Problem (1). Moreover, integrating this criterion does not increase the overall computational complexity of Algorithm 1, since the norm of $\widehat{\mathcal{G}}F(x_{t},y_{t}^{K_{t}},v_{t}^{N_{t}})$ is already computed at Step 20.
Similar to Proposition 2 in yang2024tuning , we derive upper bounds on the total number of iterations required to solve both the lower-level problem in Problem (1) and the linear system (4). However, owing to the geometric structures inherent in Riemannian manifolds, these bounds cannot be directly inferred from (yang2024tuning, , Proposition 2) and require a meticulous analysis.
@@
Suppose that Assumptions 3.1, 3.2, 3.3, and 3.4 hold. Then, for any $0\leq t\leq T$, the numbers of iterations $K_{t}$ and $N_{t}$ required in Algorithm 1 satisfy:
@@
Remark 3.1.
Proposition 3.2 provides upper bounds for the total iterations of the resolutions of the lower-level problem in Problem (1) and the linear system (4). Since $1/\log(1+\epsilon_{y})$ is of the same order as $1/\epsilon_{y}$, it holds that $K_{t}=\mathcal{O}(1/\epsilon_{y})$, which matches the standard AdaGrad-Norm for solving strongly convex problems xie2020linear . Additionally, we also have $N_{t}=\mathcal{O}(1/\epsilon_{v})$ and $N_{t}=\mathcal{O}(\log(1/\epsilon_{v}))$ when employing gradient descent and conjugate gradient methods, respectively. Moreover, similar to AdaGrad-Norm xie2020linear , the step size adaptation for the lower-level problem proceeds in two stages: Stage 1 requires at most $\mathcal{O}(1/\log(1+\epsilon_{y}))$ iterations, while Stage 2 requires at most $\mathcal{O}(\log(1/\epsilon_{y}))$ iterations. Furthermore, due to geometric properties of Riemannian manifolds, these bounds cannot be directly derived from yang2024tuning and introduce additional technical challenges. Specifically, the constants $C_{b},C_{c}$, $\bar{b}$, $\bar{c}$, $\tilde{b}$, and $\tilde{c}$ are all related to curvature $\zeta$, and increase as $\zeta$ increases.
@@
Theorem 3.1 demonstrates that our proposed adaptive algorithm achieves a convergence rate matching standard non-adaptive (Riemannian) bilevel algorithms (e.g., ghadimi2018approximation ; ji2021bilevel ; ji2022will ; han2024framework ; li2025riemannian ), demonstrating its computational efficiency. Moreover, although Algorithm 1 shares a similar algorithmic structure with D-TFBO yang2024tuning and some Euclidean analyses extend naturally to Riemannian settings, its convergence analysis still presents significant theoretical challenges. For example, the geometric structure of Riemannian manifolds necessitates the use of the trigonometric distance bound (Lemma B.1) instead of its Euclidean counterpart and requires incorporating the curvature constant into the step size adaptation (cf. Proposition 3.2), which substantially increases the analytical difficulty. We subsequently derive the complexity bound for Algorithm 1.
@@
Corollary 3.1 demonstrates that the gradient complexity ($G_{g}$) and Hessian-vector product complexity ($HV_{g}$) of AdaRHD-GD surpass those of non-adaptive Riemannian bilevel optimization methods han2024framework ; li2025riemannian by a factor of $1/\epsilon$. This gap originates from the additional iterations needed to guarantee the solution of the lower-level problem in Problem (1) and linear system 4, which is necessitated by the lack of prior knowledge regarding the strong convexity, Lipschitzness, and curvature constants.
Designing algorithms to eliminate the $1/\epsilon$-gap is an interesting future direction. Potential strategies to bridge this gap could be drawn from adaptive methods in Riemannian optimization, such as sharpness-aware minimization on Riemannian manifolds (Riemannian SAM) (yun2023riemannian, ), Riemannian natural gradient descent (RNGD) (hu2024riemannian, ), and the framework for Riemannian adaptive optimization (sakai2024general, ). Riemannian extensions of Euclidean approaches like adaptive accelerated gradient descent (malitsky2019adaptive, ), adaptive proximal gradient method (adaPGM) (latafat2024adaptive, ), adaptive Barzilai-Borwein algorithm (AdaBB) (zhou2025adabb, ), and adaptive Nesterov accelerated gradient (AdaNAG) (suh2025adaptive, ) may also be viable alternatives to the inverse of cumulative gradient norms strategy (xie2020linear, ) employed in Algorithm 1.
@@ 3.4 Extend to retraction mapping
All the iterates of the lower-level problem generated by Algorithm 3 lie in a bounded set belonging to ${\mathcal{M}}_{y}$ that contains the optimal solution, i.e., there exists a constant $\bar{D}>0$ such that $d(y_{t}^{k},y^{*}(x_{t}))\leq\bar{D}$ holds for all $t\geq 0$ and $k\geq 0$.
@@
Given $z_{1}\in{\mathcal{U}}\subseteq{\mathcal{M}}$ (here $\mathcal{M}$} can be ${\mathcal{M}}_{x}$ or ${\mathcal{M}}_{y}$) and $u\in T_{x}{\mathcal{M}}$, let $z_{2}={\rm Retr}_{z_{1}}(u)$. There exist constants $c_{u}\geq 1,c_{R}\geq 0$ such that $d^{2}(z_{1},z_{2})\leq c_{u}\|u\|_{z_{1}}^{2}$ and $\|{\rm Exp}^{-1}_{z_{1}}(z_{2})-u\|_{z_{1}}\leq c_{R}\|u\|_{z_{1}}^{2}$.
Assumption 3.6 is standard (e.g., (han2024framework, , Assumption 2) and (kasai2018riemannian, , Assumption 1)) in bounding the error between the exponential mapping and retraction mapping, given that retraction mapping is a first-order approximation to the exponential mapping (han2024framework, ).
@@
Suppose that Assumptions 3.1, 3.2, 3.3, 3.4, 3.5, and 3.6 hold. The sequence $\{x_{t}\}_{t=0}^{T}$ generated by Algorithm 3 satisfies,
@@
Algorithm 3 achieves a convergence rate nearly identical to that of Algorithm 1 while being more practical to implement. Furthermore, the complexity analysis of Algorithm 3 closely aligns with that of Algorithm 1 and is, therefore, omitted here.
@@ 4 Experimental results
In this section, adhering to the experimental framework established by han2024framework , we conduct comprehensive experiments to evaluate our algorithm against the RHGD method. Since the RieBO algorithm proposed by li2025riemannian shares the same algorithmic framework as RHGD, we categorize them as a single method for comparison. We designate RHGD-50 and RHGD-20 as configurations with maximum iteration limits for solving the lower-level problem in RHGD set to 50 and 20, respectively. For consistency, we exclusively employ Algorithm 3 in this section due to its computational efficiency in practice. Detailed experimental settings and additional experiments are provided in Appendix I, and our codes are available at https://github.com/RufengXiao/AdaRHD.